# C3D: Cascade Control with Change Point Detection and Deep Koopman Learning for Autonomous Surface Vehicles

### v1

@@ I INTRODUCTION
Extensive research efforts have focused on the study of the station keeping, for instance, developing an ASV for autonomous missions in challenging surf zone areas, utilizing stereo vision to counter wave disturbance [7]. Further studies have explored differential drive ASVs with azimuth propulsion and different control techniques such as PD, backstepping, and sliding mode controllers, including wind disturbance compensation [5, 6]. Other work has centered its attention on addressing the station keeping problem for underactuated vehicles under moderate wind conditions with relative success [8]. In more recent studies, simulations have been used to compare the implementation of a nonlinear controller for station keeping and highlight its advantages over a reinforcement learning controller [sinisterra2020nonlinear], while others have utilized simulations to assess the performance of different controllers for station keeping on underactuated ASVs, focusing on the energy efficiency [oh2020comparison]. Other work has been proposed using a cascaded controller for surge, sway, and yaw control, reducing actuation energy by computing the average disturbance direction [4].Finally, a data-driven Model Predictive Control (MPC) method is proposed to model the nonlinear ASV systems under wind disturbances and perform station keeping with an underactuated vehicle. All of this work fails to consider the influence of change in system dynamics online.
@@ III-A Validation in Simulation
To evaluate the control strategy for time-varying systems, 4 tests were carried out in the Gazebo simulator using different dynamics changes. The simulation validation is done on the desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU. The first test has no system changes. The purpose of this test is to verify the performance of the baseline methods when the vehicle is operating under its nominal conditions, as shown in 6(a). For the second test, the change is decreasing the right motor efficiency from 1 to 0.8. As for the third test, the changes are changing the angle of the right rudder by 1818$\circ$18 ∘ plus the decreased right motor efficiency. This will cause one motor to generate lateral force and make the vehicle deviate from the planned path. For the last test, a 100 kg load is added to the vehicle in addition to the right trim and right motor efficiency. This alters the mass and inertia of the vehicle and decreases its mobility.
@@ III-B Experimental Validation
To evaluate the control framework, several tests were defined to observe its performance. The first test is to test the nominal system as shown in 1 without any changes. For the second test, the change is decreasing the left motor efficiency from 1 to $0.8$. As for the third test, the changes are tilting the angle of the right engine by $30^{\circ}$ plus the decreased right motor efficiency shown in Fig. 6(d). This will cause one pontoon to be more submerged and generate more drag as shown in Fig. 6(e). For the last test, on top of the right trim and right motor efficiency, a $23$ kg sandbag is added on the right side of the vehicle as shown in 6(f), which changes the mass and inertia of the vehicle and makes it more difficult to rotate.

### v3

@@ I INTRODUCTION
Extensive research efforts have focused on the study of the station keeping, for instance, developing an ASV for autonomous missions in challenging surf zone areas, utilizing stereo vision to counter wave disturbance [7]. Further studies have explored differential drive ASVs with azimuth propulsion and different control techniques such as PD, backstepping, and sliding mode controllers, including wind disturbance compensation [5, 6]. Other work has centered its attention on addressing the station keeping problem for underactuated vehicles under moderate wind conditions with relative success [8]. In more recent studies, simulations have been used to compare the implementation of a nonlinear controller for station keeping and highlight its advantages over a reinforcement learning controller [9], while others have utilized simulations to assess the performance of different controllers for station keeping on underactuated ASVs, focusing on the energy efficiency [10]. Other work has been proposed using a cascaded controller for surge, sway, and yaw control, reducing actuation energy by computing the average disturbance direction [4]. Finally, a data-driven Model Predictive Control (MPC) method is proposed to model the nonlinear ASV systems under wind disturbances and perform station keeping with an underactuated vehicle [11]. All of this work fails to consider the influence of change in system dynamics online.
@@ III-A Validation in Simulation
To evaluate the control strategy for time-varying systems, 4 tests were carried out in the Gazebo simulator using different dynamics changes. The simulation validation is done on the desktop computer with a GeForce 2080 GPU and an Intel Core i7-8700 CPU. The first test has no system changes. The purpose of this test is to verify the performance of the baseline methods when the vehicle is operating under its nominal conditions, as shown in 7(a). For the second test, the change is decreasing the right motor efficiency from 1 to 0.8. As for the third test, the changes are changing the angle of the right rudder by 1818$\circ$18 ∘ plus the decreased right motor efficiency. This will cause one motor to generate lateral force and make the vehicle deviate from the planned path. For the last test, a 100 kg load is added to the vehicle in addition to the right trim and right motor efficiency. This alters the mass and inertia of the vehicle and decreases its mobility.
@@ III-B Experimental Validation
To evaluate the control framework, several tests were defined to observe its performance. The first test is to test the nominal system as shown in 1 without any changes. For the second test, the change is decreasing the left motor efficiency from 1 to $0.8$. As for the third test, the changes are tilting the angle of the right engine by $30^{\circ}$ plus the decreased right motor efficiency shown in Fig. 7(d). This will cause one pontoon to be more submerged and generate more drag as shown in Fig. 7(e). For the last test, on top of the right trim and right motor efficiency, a $23$ kg sandbag is added on the right side of the vehicle as shown in 7(f), which changes the mass and inertia of the vehicle and makes it more difficult to rotate.